OpenStudy (anonymous):

Which equation represents a line through (3, 5) that is perpendicular to y = 2x - 5. Answer 2y = -x + 13 2y - x = 13 2y = x + 13 2y + x = -13

Parth (parthkohli):

First, find the negative reciprocal of the slope. What do you get?

Parth (parthkohli):

What is the negative reciprocal of 2?

Parth (parthkohli):

At least, you can tell me what the reciprocal of 2 is.

OpenStudy (anonymous):

A reciprocal of x is 1/x and the negative reciprocal is -1/x

OpenStudy (anonymous):

Then, use the point-slope formula of (y-ysub1)/(x-xsub1) = (negative reciprocal of slope of original equation).

OpenStudy (anonymous):

ysub1 and xsub1 are specific values of y and x and is the point (3,5).

OpenStudy (anonymous):

Yes, I could give you the answer, but then I will get my wrists slapped here. The best I'm allowed to do is walk you through it. Where are you at with your understanding to the problem at this moment?

hartnn (hartnn):

compare your equation with y=mx+c, slope = m so u get slope of your line as 2. now product of slopes of perpendicular lines is -1 hence the slope of required line will be -1/2 so u have equation of required line as y=(-1/2)x +c to find c , use (x,y)=(3,5) so 5=(-1/2)*3+c c=5+3/2=13/2 so equation : y=(-1/2)x+13/2 or 2y = -x+13 did u understand this ?

OpenStudy (anonymous):

you couldnt have explained it any better THANKS SO MUCH:)

hartnn (hartnn):

welcome :)